# Effective Learning with Node Perturbation in Multi-Layer Neural Networks

### v4

@@ 3.4 Node perturbation allows learning in noisy systems
Figure 4: Performance of DANP when applied with one clean network pass and one noisy network pass versus with two noisy network passes. Results are reported for a three-layer fully-connected network trained for CIFAR-10 classification (comparable to Figure 3, Top panel). Curves report mean train and test accuracy. Error bars indicate the maximal and minimal accuracy obtained for five random seeds.
Figure 4 shows that computing updates based on a set of two noisy passes, rather than a clean and noisy pass, produces extremely similar learning dynamics for DANP with some minimal loss in performance. DNP, in contrast, learns more slowly during early training and becomes unstable late in training. It also does not reach the same level of test performance as it does without the noisy baseline. The similarity of the speed of learning and performance levels for DANP suggests that clean network passes may provide little additional benefit for the computation of updates in this regime. These results are most promising for application of DANP to systems in which noise is an inherent property and cannot be selectively switched off or read out.
@@ 4 Discussion
In this work, we explored several formulations of NP learning and also introduced a layer-wise decorrelation method which strongly outperforms the baseline implementations. In our results we show robust speedups in training across architectures using the developed learning rules compared to regular NP. This speedup is sufficient to suggest that our alternative formulations of NP could prove to be competitive with traditional BP in certain contexts. We attribute this increased efficacy to a more efficient credit assignment by virtue of decorrelated input features at every layer, as well as an attenuation of the bias in NP’s updates caused by their covariance.
The performance of all three methods is substantially improved by including decorrelation. The inclusion of decorrelation does appear to have an impact on the level of overfitting in these systems, with our NP-trained networks having a somewhat lower test accuracy compared to BP. This fits with recent research which suggests an overfitting impact of data whitening (Wadia et al., 2021). In this respect, further investigation is warranted.
When comparing NP, INP and ANP, a number of observations are in order. First, INP yields the best results stemming from its robust relationship to the directional derivative and underlying gradient. The INP method does, however, require running as many forward passes as there are layers in the network (plus one ‘clean’ pass). This is a drawback for this method and should be considered when analysing its performance, though it should also be considered that it could remain efficient on hardware in which all forward passes are parallelizable. Second, decorrelated variants of NP and ANP both perform well. An additional advantage of the latter though is that one does not need explicit access to the noise vectors, which has important implications for the development of physical learning machines, as described below.
Noise-based learning approaches might also be ideally suited for implementation on neuromorphic hardware (Kaspar et al., 2021). First, as demonstrated in Figure 4, our proposed DANP algorithm scales well even when there is no access to a ‘clean’ model pass. Furthermore, DANP does not require access to the noise signal itself, in contrast to NP or INP. This means that such an approach could be ideally suited for implementation in noisy physical devices (Gokmen, 2021) even in case the noise cannot be measured. Even on traditional hardware architectures, forward passes are often easier to optimize than backward passes and are often significantly faster to compute. This can be especially true for neuromorphic computing approaches, where backward passes require automatic differentiation implementations and a separate computational pipeline for backward passes (Zenke and Neftci, 2021). In these cases, a noise-based approach to learning could prove highly efficient.
@@
Though the presented methods for effective noise-based learning show a great deal of promise, there are a number of additional research steps to be taken. The architectures considered are relatively shallow, and thus an investigation into how well this approach scales for very deep networks would be beneficial. Testing the scalability of these approaches to tasks of greater complexity is also crucial, as are their application to other network architectures such as residual networks, recurrent networks and attention-based models.
@@ Appendix A Weights propagate noise correlations
Traditional NP updates have a number of sources of error. One reason is that use of the injected noise $\mathbf{\epsilon}_{l}$ to compute the learning signal ignores how the output activity of layer $l$ is not only impacted by the noise injected directly into it, but also by the cumulative effect of perturbations added to previous layers. If the noise is randomly sampled from a Gaussian distribution with mean zero, one might be tempted to simply assume that the perturbations from previous layers cancel out in expectation, but this assumption ignores the correlations introduced into these perturbations by the network’s weight matrices at all preceding layers. Multiplying a random vector by a non-orthogonal matrix will introduce a non-random covariance structure into it.
To see why, consider a transformation
$\mathbf{W}\mathbf{x}=\mathbf{y}\,,$ (7)
where $W$ is a randomly initialized weight matrix and $x$ is an uncorrelated input vector for which the following holds in expectation:
$\langle\mathbf{x}\mathbf{x}^{\top}\rangle=\mathbf{I}$ (8)
with $\mathbf{I}$ the identity matrix. The covariance matrix of $\mathbf{y}$ can be described as
@@ Appendix B The directional derivative as a measure of the gradient
$\mathbf{g}_{l}=N_{l}\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}\,,$
@@
$\mathbf{g}$ $=N\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}^{\top}\,$
$=N\left\langle\nabla\mathcal{L}^{\top}\frac{\mathbf{v}}{||\mathbf{v}||}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}^{\top}$
$=N\left(\nabla\mathcal{L}^{\top}\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle_{\mathbf{v}}\right)^{\top}$
$=N\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle_{\mathbf{v}}^{\top}\nabla\mathcal{L}$
$=N\frac{\mathbf{I}}{N}\nabla\mathcal{L}$
$=\nabla\mathcal{L}\,.$
Note that in the second to last step of the above derivation, we have taken the definition of our noise distribution as being composed of independent noise with a given variance, $\mathbf{\Sigma}=\sigma^{2}\mathbf{I}$ and combined this with the impact of normalizing every vector. Normalizing every vector means that the correlations $\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle=\left\langle\frac{\mathbf{v}}{||\mathbf{v}||}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle$ are not changed in sign but only in scale, as the noise vectors are now all scaled to lie on a unit sphere (without any rotation). In our specific case of a diagonal correlation matrix, this expectation is also therefore diagonal and all diagonal elements are equal. The average value of these diagonal elements is then simply the variance ${1}/{N}$ of a random unit vector in $N$ dimensional space. For other noise distributions, in which there does exist cross-correlation between the noise elements, this is no longer an appropriate treatment. Instead, one would have to multiply by the inverse of the correlation matrix to fully recover the gradient values.
@@ Appendix C Derivation of activity-based node perturbation
This rule has been derived in the main text in order to optimally make use of noise to determine the directional derivative with respect to a single layer of a deep network, $l$, while all other layers receive no noise. For this purpose, the noise vector is defined such that $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise.
In moving from INP to ANP, we aim to fulfil a number of conditions:
@@
These goals can be accomplished in a simple manner. To achieve the first goal, we can simply treat the whole network as if it is a single layer (even if it is not). The only change required for this modification is to assume that this rule holds even if $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ for $k\in[1,\ldots,L]$. I.e. that noise is injected for all layers and all layers are simultaneously updated.
To achieve the second goal, we can assume that one does not have access to the noise vector directly, but instead has access to the output activations from a clean and noisy pass, $\mathbf{a}_{l}$ and $\mathbf{\tilde{a}}_{l}$ respectively. Thus, rather than measuring the noise directly, one can measure the impact of all noise upon the network by substituting the noise vector, $\mathbf{v}$ for the activity difference $\delta\mathbf{a}=\mathbf{\tilde{a}}-\mathbf{a}$. Thus, with these two modifications we arrive at the ANP learning rule
$\Delta\mathbf{W}_{l}^{\textrm{ANP}}=N\>\>\delta{\mathcal{L}}\frac{\delta{\mathbf{a}_{l}}}{||\delta\mathbf{{a}}||^{2}}\>\mathbf{x}_{l-1}^{\top}\,,$
@@
Appendix D Algorithm pseudocode

### v5

@@ 3.4 Node perturbation allows learning in noisy systems
Figure 4: Performance of DNP and DANP when using two noisy network passes. Results are reported for a three-layer fully-connected network trained for CIFAR-10 classification (comparable to Figure 3, bottom panel). Curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds.
Figure 4 shows that computing updates based on a set of two noisy passes, rather than a clean and noisy pass, produces extremely similar learning dynamics for DANP with some minimal loss in performance. DNP, in contrast, learns more slowly during early training and becomes unstable late in training, decreasing in train accuracy. It also does not reach the same level of test performance as it does without the noisy baseline. The similarity of the speed of learning and performance levels for DANP suggests that clean network passes may provide little additional benefit for the computation of updates in this regime. These results are most promising for application of DANP to systems in which noise is an inherent property and cannot be selectively switched off or read out.
3.5 Resampling allows for scaling to larger systems
To further investigate the capabilities of our approach, we apply BP, DNP, DANP and DINP to the more challenging Tiny ImageNet dataset Le & Yang (2015) using a deeper architecture consisting of three convolutional layers and three fully connected layers. The Tiny ImageNet dataset has 200 classes and thus has a larger output space than CIFAR-10. Larger output spaces are known to be challenging for traditional NP-style algorithms (Hiratani et al., 2022). RGB-values for the images, which are ImageNet images downsized to $64\times 64$ pixels, were standardized as a preprocessing step. For network architecture and learning rates, see Appendix F.
Figure 5: Train and test accuracy for BP, DNP, DANP and DINP on Tiny ImageNet. DANP and DINP applied with 1, 10, 30 and 100 noise samples per update step. DNP applied with 100 noise samples per update step.
Figure 5 shows that, in their naive implementation, DNP, DANP and DINP do not appreciably learn this task. We demonstrate that, for such a difficult task, this is due to the accuracy of the gradient estimation. To show this, we resampled the random noise 10, 30 and 100 times before applying each update, effectively carrying out 10/30/100 noisy passes to compare against each single clean pass. For DNP, only an experiment with 100 iterations was included, as this algorithm did not match the performance of DINP and DANP. Results again demonstrate that both DANP and DINP improve meaningfully on DNP in deeper networks. The additional noise sampling comes at significant computational expense but could in principle be parallelized with sufficiently powerful hardware. The results indeed show that, as our theory suggests, sampling the noise distribution multiple times per update step significantly improves alignment with backpropagation, with DINP with 100 samples per update closely approximating BP’s learning trajectory. See Appendix I, Table LABEL:tab:peakacc_tin for peak accuracies.
@@ 4 Discussion
In this work, we explored several formulations of NP learning and also introduced a layer-wise decorrelation method which strongly outperforms baseline implementations. In our results we show robust speedups in training compared to regular NP, suggesting that our alternative formulations of NP could prove to be competitive with traditional BP in certain contexts. We attribute this increased efficacy to a more efficient credit assignment by virtue of decorrelated input features at every layer, as well as an attenuation of the bias in NP updates caused by their covariance.
The performance of all of our proposed methods is substantially improved by including decorrelation.
When comparing NP, INP and ANP, a number of observations are in order. First, INP yields the best results stemming from its robust relationship to the directional derivative and underlying gradient. The INP method does, however, require running as many forward passes as there are layers in the network (plus one ‘clean’ pass). This is a drawback for this method and should be considered when analyzing its performance, though it should also be considered that it could remain efficient on hardware in which all forward passes are parallelizable. Second, decorrelated variants of NP and ANP both perform well.
Noise-based learning approaches might also be ideally suited for implementation on neuromorphic hardware (Kaspar et al., 2021). First, as demonstrated in Figure 4, our proposed DANP algorithm scales well even when there is no access to a ‘clean’ model pass. Furthermore, DANP does not require access to the noise signal itself, in contrast to NP. This means that such an approach could be ideally suited for implementation in noisy physical devices (Gokmen, 2021) even in case the noise cannot be measured. Even on traditional hardware architectures, forward passes are often easier to optimize than backward passes and are often significantly faster to compute. This can be especially true for neuromorphic computing approaches, where backward passes require automatic differentiation implementations and a separate computational pipeline for backward passes (Zenke & Neftci, 2021). In these cases, a noise-based approach to learning could prove highly efficient.
@@
Though the presented methods for effective noise-based learning show a great deal of promise, there are a number of additional research steps to be taken. The architectures considered are still relatively shallow, and thus an investigation into how well this approach scales for very deep networks would be beneficial. Testing the scalability of these approaches to tasks of greater complexity is also crucial, as are their application to other network architectures such as residual networks, recurrent networks and attention-based models. Our scaling results do demonstrate that by resampling the noise we achieve better and better alignment with the gradient direction, by virtue of the formal correspondence between directional derivatives and our formulations of node perturbation. In this sense, our approach can be viewed as an alternative to forward accumulation, where we exchange memory complexity (maintaining all partial derivatives) for time complexity (repeated forward passes with resampled noise vectors).
@@ Appendix A Weights propagate noise correlations
Traditional NP updates have a number of sources of error. One reason is that use of the injected noise $\mathbf{\epsilon}_{l}$ to compute the learning signal ignores how the output activity of layer $l$ is not only impacted by the noise injected directly into it, but also by the cumulative effect of perturbations added to previous layers. If the noise is randomly sampled from a Gaussian distribution with mean zero, one might be tempted to simply assume that the perturbations from previous layers cancel out in expectation, but this assumption ignores the correlations introduced into these perturbations by the network’s weight matrices at all preceding layers. Multiplying a random vector by a non-orthogonal matrix will introduce a non-random covariance structure into it. To see why, consider a transformation $\mathbf{W}\mathbf{x}=\mathbf{y}$, where $\mathbf{W}$ is a randomly initialized weight matrix and $\mathbf{x}$ is an uncorrelated input vector for which it holds that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle=\mathbf{I}$ with $\mathbf{I}$ the identity matrix. The covariance matrix of $\mathbf{y}$ can be described as
@@ Appendix B The directional derivative as a measure of the gradient
$\mathbf{g}_{l}=N_{l}\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}$
@@
$\mathbf{g}$ $=N\left\langle{\nabla}_{\mathbf{v}}\mathcal{L}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}^{\top}=N\left\langle\nabla\mathcal{L}^{\top}\frac{\mathbf{v}}{||\mathbf{v}||}\frac{\mathbf{v}^{\top}}{||\mathbf{v}||}\right\rangle_{\mathbf{v}}^{\top}=N\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle_{\mathbf{v}}^{\top}\nabla\mathcal{L}\,.$
If we assume that our noise distribution is composed of independent noise with a given variance, $\mathbf{\Sigma}=\sigma^{2}\mathbf{I}$ then we obtain
$\mathbf{g}$ $=N\left\langle\frac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle_{\mathbf{v}}^{\top}\nabla\mathcal{L}\propto N\frac{\mathbf{I}}{N}\nabla\mathcal{L}=\nabla\mathcal{L}\,.$
Note that normalization of every vector means that the correlations $\left\langle\ifrac{\mathbf{v}\mathbf{v}^{\top}}{||\mathbf{v}||^{2}}\right\rangle$ are not changed in sign but only in scale, as the noise vectors are now all scaled to lie on a unit sphere (without any rotation). In our specific case of a diagonal correlation matrix, this expectation is also therefore diagonal and all diagonal elements are equal. The average value of these diagonal elements is then simply the variance ${1}/{N}$ of a random unit vector in $N$ dimensional space. For other noise distributions, in which there does exist cross-correlation between the noise elements, this is no longer an appropriate treatment. Instead, one would have to multiply by the inverse of the correlation matrix to fully recover the gradient values.
@@ Appendix C Derivation of activity-based node perturbation
This rule has been derived in the main text in order to optimally make use of noise to determine the directional derivative with respect to a single layer of a deep network, $l$, while all other layers receive no noise. For this purpose, the noise vector is defined such that $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. In moving from INP to ANP, we aim to fulfill a number of conditions:
@@
These goals can be accomplished in a simple manner. To achieve the first goal, we can simply treat the whole network as if it is a single layer (even if it is not). The only change required for this modification is to assume that this rule holds even if $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ for $k\in[1,\ldots,L]$. I.e., that noise is injected for all layers and all layers are simultaneously updated. To achieve the second goal, we can assume that one does not have access to the noise vector directly, but instead has access to the output activations from a clean and noisy pass, $\mathbf{a}_{l}$ and $\mathbf{\tilde{a}}_{l}$ respectively. Thus, rather than measuring the noise directly, one can measure the impact of all noise upon the network by substituting the noise vector, $\mathbf{v}$ for the activity difference $\delta\mathbf{a}=\mathbf{\tilde{a}}-\mathbf{a}$. Thus, with these two modifications we arrive at the ANP learning rule
$\Delta\mathbf{W}_{l}^{\textrm{ANP}}=N\>\delta{\mathcal{L}}\frac{\delta{\mathbf{a}_{l}}}{||\delta\mathbf{{a}}||^{2}}\>\mathbf{x}_{l-1}^{\top}$
@@
Appendix D Bias in NP compared to ANP
In this section, we analyze the deviation of NP’s and ANP’s updates from the true gradient in a simple two-layer neural network, focusing on the effects of noise propagation through the network’s layers. Unlike existing work in the literature, we do not assume infinitely small perturbations, or a linear network, to provide a quantification of bias in NP’s and ANP’s updates. In practice, perturbations applied to networks cannot be infinitely small and are even limited by precision of floating point representations to be larger than some minimum size.
Consider a two-layer neural network with weight matrices $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, no biases, and a nonlinear activation function $f(\cdot)$ applied after the first layer. The network’s output, given an input $x$, is expressed as:
$\mathbf{y}_{2}=\mathbf{W}_{2}f(\mathbf{W}_{1}\mathbf{x})\,.$
Noise is introduced into the network as small perturbations $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ at the outputs of the first and second layers, respectively. With this perturbation, the perturbed output becomes:
$\tilde{\mathbf{y}}_{2}=\mathbf{W}_{2}\big{(}f(\mathbf{W}_{1}\mathbf{x}+\mathbf{v}_{1})+\mathbf{v}_{2}\big{)}\,.$
Given an arbitrary loss function $L$, the loss differential is given by $L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2})$. NP’s updates are based on a product between this loss differential and the noise. Examining this for matrix $\mathbf{W}_{2}$, this corresponds to a product between the noise differential and vector $\mathbf{v}_{2}$. To fully show the impact of this update, we expand the loss differential by a Taylor expansion beyond the first term (i.e. assuming that our loss difference is not infinitely small or linearizable) and then further expand the activity difference by Taylor expansion. As such,
$\langle\Delta\mathbf{W}_{2}^{\text{NP}}\rangle$ $=\frac{1}{h}\left\langle(L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2}))\cdot\mathbf{v}_{2}\right\rangle f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\Bigg{\langle}\mathbf{v}_{2}\Bigg{[}\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{2!}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla^{2}_{\mathbf{y}_{2}}L\;(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{3!}\cdots\Bigg{]}^{\top}\Bigg{\rangle}f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\left\langle\mathbf{v}_{2}\left[\nabla_{\mathbf{y}_{2}}L^{\top}\left(\mathbf{v}_{2}+\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}^{\top}\mathbf{v}_{1}+\frac{1}{2!}\mathbf{v}_{1}^{\top}\nabla^{2}_{\mathbf{y}_{1}}\mathbf{y}_{2}\mathbf{v}_{1}+\frac{1}{3!}\cdots\right)\cdots\right]^{\top}\right\rangle f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\left\langle\mathbf{v}_{2}\left[\left(\mathbf{v}_{2}+\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}^{\top}\mathbf{v}_{1}+\frac{1}{2!}\mathbf{v}_{1}^{\top}\nabla^{2}_{\mathbf{y}_{1}}\mathbf{y}_{2}\mathbf{v}_{1}+\frac{1}{3!}\cdots\right)^{\top}\nabla_{\mathbf{y}_{2}}L\cdots\right]\right\rangle f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\bigg{\langle}\mathbf{v}_{2}\mathbf{v}_{2}^{\top}+\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}+\frac{1}{2!}\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\nabla^{2}_{\mathbf{y}_{1}}\mathbf{y}_{2}^{\top}\mathbf{v}_{1}+\frac{1}{3!}\cdots\bigg{\rangle}\nabla_{\mathbf{y}_{2}}L\;f(\mathbf{y}_{1})^{\top}$
=$\cdots$= ⋯
The above Taylor expansion demonstrates that above the linear terms ($\mathbf{v}_{2}\mathbf{v}_{2}^{\top}+\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}\approx\mathbf{I}+\mathbf{0}$, in the ideal noise case), there are terms involving the correlation between noise vector $\mathbf{v}_{2}$ and higher order terms invoving $\mathbf{v}_{1}$, beginning with the Hessian but also including third, fourth and higher derivative terms. We cannot, in general, say anything about the unbiasedness of these terms and these are sources of unexplained error in existing work.
By comparison, ANP’s Taylor expansion has a much more interpretable correlation structure. Again looking at weight updates for layer two:
$\langle\Delta\mathbf{W}_{2}^{\text{ANP}}\rangle$ $=\frac{1}{h}\Big{\langle}\big{(}L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y})\big{)}\cdot(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{[}\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{2!}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla^{2}_{\mathbf{y}_{2}}L(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{3!}\dots\Big{]}^{\top}\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla_{\mathbf{y}_{2}}L+\frac{1}{2!}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla^{2}_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{3!}\dots\Big{]}^{\top}\Big{\rangle}f(\mathbf{y}_{1})^{\top}\,.$
This simplifies all assumptions regarding higher-order noise propagation terms, and reduces the algorithms bias down to the correlation in overall activity differences ($\langle(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\rangle$). This is a much more interpretable bias term which could even be targeted directly to be made diagonal if a practitioner was interested in doing so.
Appendix E Algorithm pseudocode